OpenStudy (candyme):

Use addition to solve this system of equations. Please show how to do this because I want to know how. 8x+y=1 -8x-4y=3

OpenStudy (anonymous):

Add the two together: -3y=4 y=-4/3

OpenStudy (saifoo.khan):

x = 7/24 y = -4/3